Assessment

Solve the following.

1. The perimeter of a rectangle is 24 cm. Express the area of the rectangle in terms of the width x.

2. The area of a rectangle is 85 cm. Express the perimeter of the rectangle as a function of the width x.

3. A piece of wire y cm long is bent into a circle. Express the area of the circle as a function of the circumference y.

4.  Express the length of the radius of a circle as a function of the 

5. Express the length of the diameter of a circle as a function of the circumference.

6.The perimeter of a rectangle is 50 centimeters. Express its area A as a function of the length x.

7. A wire of length x is bent into the shape of a square. Express the area of the square as a function of x. 


1
Expert's answer
2021-08-25T05:36:45-0400

1."Width" "=" "x"

"Perimeter=2(L+W)"

"Length=12-x"

Area of a rectangle is represented as: "L\times W"

"Area=x(12-x)"


"Area=12x-x^2"


2."Width=x"

"Area=L\times W"

"Length=\frac{85}{x}"

"Perimeter=2(\frac{85}{x}+x)"


"Perimeter=\frac{170}{x}+2x"


3."Circumference=2\pi r"

"y=2\pi r"

"r=\frac{y}{2\pi}"

"Area=\pi r^2"

"Area=\pi \times(\frac{y}{2\pi})^2"

"Area=\pi \times(\frac{y^2}{4\pi^2})"


"Area=\frac{y^2}{4\pi}"

4."Circumference=2\pi r"

"y=2\pi r"


"r=\frac{y}{2\pi}"


5. "r=\frac{y}{2\pi}"

"diameter=r+r=2r"

"2r=2(\frac{y}{2\pi})"


"2r=\frac{y}{\pi}"


6."Length=x"

"Width=25-x"

"A=L\times W"

"A=x(25-x)"


"A=25x-x^2"


7."Perimeter" "of" "a" "square=x"

Since all sides of a square are equal:

"Length" "of" "a" "sqaure=\frac{x}{4}"

"Area=Side\times Side=S\times S"

"Area=\frac{x}{4} \times \frac{x}{4}"


"Area=\frac{x^2}{16}"